Please help me solve the following Linear Equations for the unknown:
2 + 𝑥 = 1
4𝑐 – 2 = 3 − 6𝑐

Answer by Cromlix(4381)

Hi there,
2 + x = 1
x = 1 - 2
x = -1
.......
4c - 2 = 3 - 6c
4c + 6c = 3 + 2
10c = 5
c = 5/10
c = 1/2
Hope this helps :-)